Using the income statement for Times Mirror and Glass Co., compute the following ratios TIMES MIRROR AND GLASS Co. Income Statement Sales Cost of goods sold Gross profit Selling and administrative expense Lease expense Operating profit Interest expense Earnings before taxes Taxes (30%) $221,000 113,000 $108,000 43,700 11,900 52,400 7,100 45,300 18,120 $27,180 Earnings after taxes Equals income befor re interest and taxes a.Compute the interest coverage ratio. (Round your answer to 2 decimal places.) Interest coverage times b.Compute the fixed charge coverage ratio. (Round your answer to 2 decimal places.) Fixed charge coverage times The total assets for this company equal $162,000. Set up the equation for the Du Pont system of ratio analysis c.Compute the profit margin ratio. (Input your answer as a percent rounded to 2 decimal places.) Profit margin d.Compute the total asset turnover ratio. (Round your answer to 2 decimal places.) Total asset turnover times e. Compute the return on assets (investment). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.) Return on assets
Answer #1

a)Interest coverage ratio=operating profit/interest expenses
=52400/7100=7.38
b)Fixed charge coverage=(operating profit+lease expense)/(interest expense+lease expense)
=(52400+11900)/(7100+11900)=3.38
c)Profit margin=earning after tax/sales
=27180/221000=12.30%
d)TOtal asset turnover=sales/total assets
=221000/162000=1.36
e)Return on Assets=(earning after tax/Assets)
=(earning after tax/sales)*(sales/Assets)
=Profit margin*total asset turnover
=12.30%*1.36=16.73%